ABC Company is not expected to pay any dividends for the next 3 years. Beginning 4 years from today, investors expect to receive a dividend of $2 per share for 3 years and then a dividend of $3 per share for each of the next 4 years. Then dividends are expected to grow at 4% per year forever. If investors require a 15% return, what is the price per share? Do not use excel. Need to understand this writing out PVIF functions.  
Answer #1

Year

Dividend

1

D1=

$0

2

D2=

$0

3

D3=

$0

4

D4=

$2

5

D5=

$2

6

D6=

$2

7

D7=

$3

8

D8

$3

9

D9=

$3

10

D10

$3

11

D11

$3.12

(3*1.04)

12

$3.24

(3.12*1.04)

13

$3.37

(3.24*1.04)

P10=Price in year 10=D11/(R-g)=Present Value in year 10 of future cash flows from year 11 to perpetuity.

R=Required Return=15%=0.15

g= 4%=0.04

P10=3.12/(0.15-0.04)=$28.36

Current Price Per Share =P0=Present Value of future cash flows

Present Value (PV) of Cash Flow:

(Cash Flow)/((1+i)^N)

i=Discount Rate=Required rate of return=15%=0.15

N=Year of Cash Flow

Cash Flows and Present value of cash flows are given below:

N

A

PV=A/(1.15^N)

Year

Cash Flow

1

D1

$0

$0

2

D2

$0

$0

3

D3

$0

$0

4

D4

$2

1.143506491

(2/(1.15^4)

5

D5

$2

0.994353471

(2/(1.15^5)

6

D6

$2

0.864655192

(2/(1.15^6)

7

D7

$3

1.12781112

(3/(1.15^7)

8

D8

$3

0.980705322

(3/(1.15^8)

9

D9

$3

0.852787236

(3/(1.15^9)

10

D10

$3

0.741554118

(3/(1.15^10)

10

P10

$28.36

7.010158266

(28.36/1.15^10)

SUM

$13.72

Price Per Share =Sum of PV of future cash flows=$13.72

Current Price per share=P0=

$13.72